HUBBLE SPACE TELESCOPE DAILY REPORT #3095 PERIOD COVERED: 0000Z (UTC) 04/15/02 - 0000Z (UTC) 04/16/02 Daily Status Report as of 106/0000Z 1.0 OBSERVATIONS SCHEDULED AND ACCOMPLISHED: 1.1 Completed Two Sets of STIS/CCD 8902 (Dark Monitor-Part 2) The Space Telescope Imaging Spectrograph (CCD) was used to monitor the darks. No problems were reported. 1.2 Completed Five Sets of ACS/WFC/HRC 8947 (Weekly Test) The Advanced Camera for Surveys (WFC and HRC) was used to perform basic tests to monitor, the read noise, the development of hot pixels and test for any source of noise in ACS CCD detectors. As described in yesterday's report and HSTAR 8607, the acquisition for the third iteration of this proposal defaulted to fine lock back-up on one FGS only, possibly affecting three observations. There were no other problems reported. 1.3 Completed Three Sets of WF/PC-2 8937 (Cycle 9 Supplemental Darks pt2/3) The WF/PC-2 was used obtain three dark frames every day to provide data for monitoring and characterizing the evolution of hot pixels. No problems were encountered. 1.4 Completed Two Sets ofACS/WFC 9020 (Preliminary ACS Sensitivity) The Advanced Camera for Surveys (WFC) was used to observe a spectrophotometric standard star through each filter of each camera to assess the sensitivity of the instrument. The star is placed at the center of the aperture, and two images are taken through each filter. There were no reported problems. 1.5 Completed STIS/CCD 8904 (Bias Monitor-Part 2) The Space Telescope Imaging Spectrograph (CCD) was used to monitor the bias in the 1x1, 1x2, 2x1, and 2x2 bin settings at gain=1, and 1x1 at gain = 4 in order to build up high-S/N superbiases and track the evolution of hot columns. The proposal completed with no anomalous activity. 1.6 Completed Two Sets of STIS/CCD 8905 (Read Noise Monitor) The Space Telescope Imaging Spectrograph (CCD) was used to measure the read noise of all the amplifiers {A, B, C, D} on the STIS CCD using pairs of bias frames. Full frame and binned observations are made in both Gain 1 and Gain 4, with binning factors of 1x1, 1x2, 2x1 and 2x2. All exposures are internals. The proposal completed with no reported anomalies. 1.7 Completed STIS/CCD 8907 (Spectroscopic Flats C10) The Space Telescope Imaging Spectrograph (CCD) was used to obtain CCD flats in the spectrographic mode. There were no reported problems. 1.8 Completed Two Sets of WF/PC-2 8939 (Cycle 10 Internal Monitor) The WF/PC-2 was used to calibrate the internal monitor, to be run weekly to monitor the health of the cameras. No problems were encountered. 1.9 Completed ACS/SBC 9008 (MAMA Initial Turn-on and Anomalous Recovery Procedure) The Advance Camera for Surveys (SBC) was used for the initial turn-on of the ACS MAMA detector and to permit recovery after an anomalous shutdown. Anomalous shutdowns can occur as a result of bright object violations which trigger the Bright Scene Detection or Software Global Monitors. Anomalous shutdowns can also occur as a result of MAMA hardware problems. The Initial MAMA turn-on/recovery from anomalous shutdown consists of three tests: a signal processing electronics check, high voltage ramp-up to an intermediate voltage, and high voltage ramp-up to the full operating voltage. During each of the two high voltage ramp-ups, diagnostics are performed during a dark ACCUM. No problems were recorded. 1.10 Completed STIS/CCD 9131 (Imaging the Host Galaxies of High Redshift Type Ia Supernovae) The Space Telescope Imaging Spectrograph (CCD) was used to complete the snapshot survey of distant galaxies of known redshift which hosted supernovae {SNe} of Type Ia found via the Supernova Cosmology Project {SCP}. No problems were seen. 1.11 Completed Two Sets of STIS/CCD 8908 (CCD Imaging Flats C10) The Space Telescope Imaging Spectrograph (CCD) was used to investigate flat-field stability over a monthly period. The proposal completed with no reported problems. 1.12 Completed STIS/CCD/MA1 9486 (What Excites LINERs: The Brilliant Case of NGC 3998) The Space Telescope Imaging Spectrograph (CCD and MA1) was used to observe LINER emission that is now known to occur in many galaxies at a variety of levels, with causes as diverse as old nuclear starbursts, cooling flows, accretion disks, and jets. The early prototypes, including NGC 3998, belong to a distinct class of AGN-like LINERs with the most luminous line emission and powerful nuclear radio emission. These LINERs have ionized gaseous disks on the scale of tens of parsecs, which have now been found to be in rotation about supermassive black holes. The disks are apparently the outer parts of energy- releasing accretion disks. Powerful new UV+optical emission line diagnostics have been developed to settle the long- standing debate over whether the gas is being photoionized or shock excited. As described in yesterday's report and HSTAR 8607, the acquisition for the third iteration of this proposal defaulted to fine lock back-up on one FGS only, possibly affecting six observations. There was no other reported anomalous activity. 1.13 Completed STIS/MA2 9573 (NUV-MAMA Daily Dark Monitor) The Space Telescope Imaging Spectrograph (MA2) was used to perform daily monitoring of the NUV MAMA detector dark noise in order to monitor the effects of thermal changes on the NUV dark rate. No problems occurred. 1.14 Completed STIS/CCD/MA1 9067 (UV Detectability of Bright Quasars in the Sloan Fields) The Space Telescope Imaging Spectrograph (CCD and MA1) was used to take MAMA spectra of approximately 30 new, bright, high-redshift quasars in each of the next three cycles. The observations completed with no reported problems. 1.15 Completed STIS/CCD/MA1 9459 (The Response of the White Dwarf in WZ Sge to the Unexpected July 2001 Superoutburst) The Space Telescope Imaging Spectrograph (CCD and MA1) was used to observe WZ Sge, the most extreme dwarf nova and one of the closest known cataclysmic variables, that has undergone a superoutburst in July 2001 after 22 years in quiescence. Because of the uniqueness of this event, two DD proposals were approved, one to observe the outburst itself, and another for us to observe the early decline phase. Here it is proposed to complete our fundamental study of the response of a dwarf nova system to an outburst by continuing our UV coverage of this most extreme outbursting system during its decline to quiescence. This decline is expected to take more than 3 yrs, with the most dramatic changes occurring in the first 2 years. This once-in-a-lifetime chance to obtain high quality, high time {and spectral} resolution FUV data as the decline progresses into the critical transition from the disk- dominated phase to the bare white dwarf, provides an unique opportunity to study the response of the emerging white dwarf, whose chemical abundances, rotation and temperature variation with time bear the imprint of this extraordinary gigantic accretion event. The proposal completed with no reported problems. 1.16 Completed STIS/MA1/MA2 9120 (Planetary Nebulae In The LMC: A Study On Stellar Evolution And Populations) The Space Telescope Imaging Spectrograph (MA1 and MA2) was used to investigate the final phase of the evolution of low- and intermediate-mass stars, the Planetary Nebula {PN} ejection that is thought to provide the main source of carbon and nitrogen enrichment in galaxies. The observations completed nominally with no reported problems. 1.17 Completed STIS/MA2 8920 (Cycle 10 MAMA Dark Measurements) The Space Telescope Imaging Spectrograph (MA2) was used to perform the routine monitoring of the MAMA detector dark noise, and is the primary means of checking on health of the MAMA detectors systems through frequent monitoring of the background count rate. The proposal completed with no reported anomalous activity. 1.18 Completed STIS/CCD 9077 (Survey of the LMC Planetary Nebulae) The Space Telescope Imaging Spectrograph (CCD) was used to perform a snapshot survey of all known LMC planetary nebulae {PNe} in order to study the co-evolution of the nebulae and their central stars, and to probe the chemical enrichment history of the LMC. There were no reported problems. 1.19 Completed FGS/1 9034 (The Masses and Luminosities of Population II Stars) Fine Guidance Sensor #1R was used to observe the mass-luminosity relation {MLR} of Population II stars of which very little is currently known. With the advent of the Hipparcos Catalogue, improved distances to many spectroscopic binaries known to be Pop II systems are now available. After surveying the literature and making reasonable estimates of the secondary masses, we find 13 systems whose minimum separation should be larger than the resolution limit of FGS #1R. The observations completed nominally. 1.20 Completed WF/PC-2/STIS/CCD 9174 (Using Optically Faint Radio Sources to Pinpoint Dusty Proto-Galaxies) The WF/PC-2 and the Space Telescope Imaging Spectrograph (CCD) were used to observe a number of high redshift, dusty starburst galaxies that remain invisible in ground based observations in the Hubble Deep Field. These galaxies appear as faint radio sources with bright submillimeter counterparts, demonstrating they are very dusty systems. The luminosities of these galaxies suggest they are in the process of converting the bulk of their gas mass into stars, and will likely evolve into present day massive ellipticals. The observations completed normally, with no reported problems. 1.21 Completed STIS/CCD 9070 (A Census of Nuclear Star Clusters in Late-Type Spiral Galaxies: II. Spectroscopy and Stellar Populations) The Space Telescope Imaging Spectrograph (CCD) was used to investigate spiral galaxies that have a prominent star cluster in their dynamical center. Statistics for cluster frequency, size, and luminosity remain incomplete. The proposal completed as planned. 2.0 FLIGHT OPERATIONS SUMMARY: 2.1 Guide Star Acquisitions: Scheduled Acquisitions: 12 Successful: 12 Scheduled Re-acquisitions: 5 Successful: 5 2.2 FHST Updates: Scheduled: 24 Successful: 22 Per HSTAR 8608, the full maneuver updates scheduled for 106/084931Z and 106/085216Z failed with error results indicating tracker #1 failed. The subsequent acquisition was successful. 2.3 Operations Notes: Operations notes were utilized to: Restore ACS flight software error limits. Perform HSTAR documentation for FHST map/update failures. Using operations requests, the following actions were executed: Clear ACS flight software error counter at 105/1343Z. Genslew for proposal 9110 - slot 1 at 105/1428Z. Clear ACS SBC event flag at 105/1543Z. Set ACS SBC event flag to prevent high voltage turn-on at 105/2013Z (BALL expressed concerns that the SBC low voltage supply had not been enabled long enough (>18 hours) to allow sufficient out-gassing prior to high voltage operations). Per ROP RD-7A, ESTR re-conditioning was performed at 105/2039Z. The ephemeris table was uplinked at 106/0127Z, using ROP DF-07A. The engineering status buffer limit for SESBSLD was updated at 106/0856Z as directed by ROP DF-18A. 3.0 SIGNIFICANT FORTHCOMING EVENTS: Battery #3 Capacity Test is scheduled to occur at approximately 107/1130Z. An intercept SMS, the intercept point being approximately 109/0354Z, is being prepared. This SMS will recover NICMOS and then begin NICMOS Servicing Mission Orbital Verification (SMOV) activities. Continuation of SMOV and the gradual resumption of normal science observations and calibrations. /CAW